. In a group of 150 students, more students are taking a math class than are taking a science class this semester. If 80 are taking a science class and 25 are not taking either math or science this semester, what is the minimum number of students who could be taking both math and science this semester?

Answer by Edwin McCravy(20067)

In a group of 150 students, more students are taking a math class than are taking a science class this semester. If 80 are taking a science class and 25 are not taking either math or science this semester, what is the minimum number of students who could be taking both math and science this semester?

There are four categories of students here:

1. Those taking math but not science.
2. Those taking science but not math.
3. Those taking both math and science.
4. Those taking neither math nor science.

1. Let x = the number of students taking math but not science.
2. Let y = the number of students taking science but not math.
3. Let z = the number of students taking both math and science.
4. Let w = the number of students taking neither math nor science.

We are asked:

>>.....what is the minimum number of students who could be 
taking both math and science this semester?.....<<

So we want to know the smallest possible value of z.

>>......In a group of 150 students,......<<

Therefore, x + y + z + w = 150 

>>......more students are taking a math class than are taking 
a science class this semester......<<

Therefore x + z > y + z.  Subtracting z from both sides,
              x > y

 
>>......If 80 are taking a science class......<<

Therefore y + z = 80

>>......and 25 are not taking either math or science this 
semester,......<< 

Therefore w = 25

So we have these four facts:

x + y + z + w = 150
x > y
y + z = 80
w = 25

Substituting w = 25 in the first equation,

x + y + z + 25 = 150
     x + y + z = 125

Substitute 80 for y + z in that:

        x + 80 = 125 
             x = 45

Substitute x = 45 in x > y

            45 > y

Solve y + z = 80 for y
          y = 80 - z

Substitute 80 - z for y in 45 > y

          45 > 80 - z

           z > 80 - 45

           z > 35

So the minimum value of z is 36, for
that is the smallest possible whole
number greater than 35.

Edwin